Mohan's father was two times as old as Mohan was two year ago.Today sum of their age is 60.Find their present age.
Answer by stanbon(75887)

Mohan's father was two times as old as Mohan was two year ago.Today sum of their age is 60.Find their present age.
------
Equations:
f = 2(m-2)
f + m = 60
-----
Substitute for "f" and solve for "m":
60-m = 2m-4
3m = 64
Mohan's age = 21 1/3 yrs
--------------------------
Solve for "f":
f + (64/3) = 60
f + (64/3) = 180/3
---
f = 116/3 = 38 2/3 (father's age now)
===============
Cheers,
Stan H.
===============